# MVBench: A Comprehensive Multi-modal Video Understanding Benchmark

### 28 Nov 2023 v1

@@ 1 Introduction
Second, guided by the task definition, we design an automatic annotation paradigm to generate multi-choice QAs for each task, by converting 11 public video benchmarks with LLMs. On one hand, it can largely reduce the cost of expensive human annotations. On the other hand, these 11 benchmarks cover various complex domains and diverse scenes, ranging from first-person to third-person perspectives, and from indoor to outdoor environments. Hence, our MVBench is a preferable choice to evaluate the general capability of MLLMs for open-world temporal understanding. More importantly, these benchmarks provide the ground truth for MVBench which guarantees evaluation fairness and accuracy, avoiding biased scoring of LLMs [81, 49].
@@ 3 MVBench
In this section, we present our MVBench in detail. We first design the temporal tasks in Tab. 1, and then automatically generate multi-choice QAs for MLLM evaluation in Fig. 2.
@@ 3.1 Temporal Task Definition
Action. (1) Action Sequence: Retrieve the events occurring before or after a specific action. (2) Action Prediction: Infer the subsequent events based on the current actions. (3) Action Antonym: Distinguish the correct action from two inversely ordered actions. (4) Fine-grained Action: Identify the accurate action from a range of similar options. (5) Unexpected Action: Detect surprising actions in videos characterized by humor, creativity, or magic. Object. (6) Object Existence: Determine the existence of a specific object during a particular event. (7) Object Interaction: Identify the object that participates in a particular event. (8) Object Shuffle: Locate the final position of an object in an occlusion game. Position. (9) Moving Direction: Ascertain the trajectory of a specific object’s movement. (10) Action Localization: Determine the time period when a certain action occurs. Scene. (11) Scene transition: Determine how the scene transitions in the video. Count. (12) Action Count: Calculate how many times a specific action has been performed. (13) Moving Count: Calculate how many objects have performed a certain action. Pose. (14) Fine-grained Pose: Identify the accurate pose category from a range of similar options. Attribute. (15) Moving Attribute: Determine the appearance of a specific moving object at a given moment. (16) State Change: Determine whether the state of a certain object changes throughout the video. Character. (17) Character Order: Determine the order in which the letters appear. Cognition. (18) Egocentric Navigation: Forecast the subsequent action, based on an agent’s current navigation instructions. (19) Episodic Reasoning: Perform reasoning on the characters, events, and objects within an episode of a TV series. (20) Counterfactual Inference: Consider what might happen if a certain event occurs.
@@ 3.2 Automatic QA Generation
With the guidance of temporal task definitions, we next collect and annotate videos for each task. Specifically, we design an automatic QA generation paradigm in Fig. 2, which efficiently converts open-sourced video annotations into multi-choice QAs for evaluating MLLMs.
@@
QA Generation. Considering that not all the annotations of selected datasets follow the multi-choice QA format, we automatically convert the video annotations into this format via LLMs. Specifically, we first use ChatGPT [54] to generate a question for each video, based on the task definition. Then, we create the corresponding answer options as follows. (1) Template-Based Construction: For most questions, we construct the option candidates directly from the ground truth annotations. For example, the candidates for the Action Antonym task contain the correct action, its opposite action, and a not-sure choice. In the case of the Moving Direction task, the option candidates consist of four directions (i.e., up, down, left, right) and the stationary state. (2) LLM-Based Generation: For the Unexpected Action task in particular, we leverage ChatGPT for converting open-ended QAs into multi-choice QA with answer options. Note that, we use the multi-choice format instead of the open-ended one, for evaluation correction and fairness. This is mainly because the open-ended answer has to be scored by LLMs or user studies, which may either introduce evaluation bias or manual intervention. Ultimately, we produce 200 multi-choice QA pairs for each temporal understanding task. More details of QA generation for all the tasks can be found in the appendix.
@@ 4.2 Progressive Multi-Modal Training
| Random | - | 27.7 | 25.0 | 25.0 | 33.3 | 25.0 | 25.0 | 33.3 | 25.0 | 33.3 | 25.0 | 25.0 | 25.0 | 33.3 | 33.3 | 33.3 | 33.3 | 25.0 | 33.3 | 25.0 | 20.0 | 30.9 |
| ImageLLM: Following [11], all models take 4 frames as input, with the output embeddings concatenated before feeding into the LLM. | ||||||||||||||||||||||
@@
| VideoLLM: All models take 16 frames as input, with the exception of VideoChatGPT, which uses 100 frames. | ||||||||||||||||||||||
@@ 6 Conclusion
Appendix A Training Hyperparameters
The hyperparameters utilized in different stages of training are listed in Tab. 10. We adopt TSN [71] sampling for all the videos as previous methods [41, 72, 40]. For both Stage1 and Stage2, we employ large-scale image and video caption data, as outlined in the main manuscript. During Stage3, we make use of diverse instruction data and incorporate LoRA modules [24] into the LLM with a rank of 16, an alpha value of 32, and a dropout rate of 0.1. We apply flash attention [12] for LLM to expedite the training process.
@@ Appendix D Results on Challenging Video QA
| Random | - | 27.7 | 25.0 | 25.0 | 33.3 | 25.0 | 25.0 | 33.3 | 25.0 | 33.3 | 25.0 | 25.0 | 25.0 | 33.3 | 33.3 | 33.3 | 33.3 | 25.0 | 33.3 | 25.0 | 20.0 | 30.9 |
| GPT-4V [55] | GPT-4 | 49.5 | 80.0 | 40.0 | 30.0 | 50.0 | 60.0 | 60.0 | 90.0 | 40.0 | 20.0 | 60.0 | 100.0 | 40.0 | 20.0 | 50.0 | 40.0 | 70.0 | 50.0 | 20.0 | 60.0 | 10.0 |
| ImageLLM: Following [11], all models take 4 frames as input, with the output embeddings concatenated before feeding into the LLM. | ||||||||||||||||||||||
| Otter-I [37] | MPT-7B | 35.0 | 50.0 | 30.0 | 20.0 | 30.0 | 40.0 | 50.0 | 50.0 | 30.0 | 30.0 | 10.0 | 40.0 | 40.0 | 20.0 | 40.0 | 40.0 | 40.0 | 30.0 | 30.0 | 40.0 | 40.0 |
| InstructBLIP [11] | Vicuna-7B | 34.0 | 40.0 | 50.0 | 40.0 | 40.0 | 40.0 | 60.0 | 40.0 | 10.0 | 20.0 | 20.0 | 70.0 | 30.0 | 20.0 | 30.0 | 50.0 | 40.0 | 20.0 | 20.0 | 30.0 | 10.0 |
| LLaVA [45] | Vicuna-7B | 34.5 | 40.0 | 30.0 | 20.0 | 30.0 | 60.0 | 50.0 | 30.0 | 20.0 | 30.0 | 30.0 | 40.0 | 20.0 | 40.0 | 30.0 | 40.0 | 30.0 | 40.0 | 20.0 | 50.0 | 40.0 |
| VideoLLM: All models take 16 frames as input, with the exception of VideoChatGPT, which uses 100 frames. | ||||||||||||||||||||||
| VideoChatGPT [49] | Vicuna-7B | 32.5 | 30.0 | 30.0 | 50.0 | 30.0 | 40.0 | 70.0 | 20.0 | 50.0 | 10.0 | 20.0 | 10.0 | 30.0 | 10.0 | 30.0 | 70.0 | 20.0 | 50.0 | 10.0 | 50.0 | 20.0 |
| VideoLLaMA [95] | Vicuna-7B | 34.0 | 30.0 | 20.0 | 50.0 | 40.0 | 20.0 | 50.0 | 40.0 | 0.0 | 30.0 | 10.0 | 70.0 | 40.0 | 30.0 | 30.0 | 50.0 | 30.0 | 50.0 | 50.0 | 40.0 | 0.0 |
| VideoChat [40] | Vicuna-7B | 36.5 | 50.0 | 10.0 | 50.0 | 20.0 | 50.0 | 60.0 | 70.0 | 40.0 | 40.0 | 40.0 | 50.0 | 30.0 | 20.0 | 40.0 | 50.0 | 20.0 | 30.0 | 10.0 | 50.0 | 0.0 |
| VideoChat2 | Vicuna-7B | 35.0 | 40.0 | 40.0 | 30.0 | 30.0 | 30.0 | 40.0 | 20.0 | 50.0 | 20.0 | 30.0 | 60.0 | 60.0 | 30.0 | 20.0 | 60.0 | 20.0 | 50.0 | 30.0 | 30.0 | 10.0 |
| VideoChat2 | Vicuna-7B | 56.5 | 60.0 | 60.0 | 90.0 | 60.0 | 60.0 | 70.0 | 80.0 | 30.0 | 20.0 | 40.0 | 100.0 | 40.0 | 40.0 | 40.0 | 60.0 | 50.0 | 40.0 | 60.0 | 60.0 | 70.0 |
@@
In Tabs. 17 and 18, we extend the evaluation of our VideoChat2 to other challenging video benchmarks i.e., NExT-QA [80], STAR [78] and TVQA [34]. Different from the previous methods [90], which provide answers by comparing the likelihood of different options, we output the options directly, following the protocol of MVBench. Our results indicate that VideoChat2 not only holds its own against current SOTA methods [73, 90] on NExT-QA but also markedly outperforms them on STAR and TVQA. This underscores the effectiveness and robustness of VideoChat2.
Appendix E Comparisons with GPT-4V
We further conduct evaluations for GPT-4V [55] in Tab. 16. Given the time constraints, we randomly sample 10 multiple-choice QAs for each task. The results indicate that GPT-4V [55] achieved satisfactory performance on our MVBench, demonstrating its considerable capacity for temporal understanding. Notably, our VideoChat2 outperforms it by increasing accuracy by 7%,
@@ Appendix E Comparisons with GPT-4V
VideoChat2
58.5
2
[…]
10cm
VideoChat
42.5
3
[…]
14cm
LLaMA-Adapter
41.5
4
[…]
14cm
InstructBLIP
40.5
5
[…]
14cm
BLIP2
40.0
6
[…]
10cm
VideoChatGPT
39.5
7
[…]
14cm
LLaVA
38.5
8
[…]
10cm
VideoLLaMA
32.5
9
[…]
14cm
mPLUG-Owl-I
31.5
10
[…]
14cm
Otter-I
28.5
11
[…]
14cm
MiniGPT-4
8.0
10cm
@@
Appendix F Leaderboards and Analyses
To facilitate a clear comparison of different MLLMs, we present the leaderboards for different tasks on MVBench in Tab. 19. Overall, our VideoChat2 achieves the highest rank across 15 tasks.
@@ Appendix F Leaderboards and Analyses
Appendix G Qualitative Results

### 21 Jan 2024 v3

@@ 1 Introduction
Second, guided by the task definition, we design an automatic annotation paradigm to generate multiple-choice QAs for each task, by converting 11 public video benchmarks with LLMs. On one hand, it can largely reduce the cost of expensive human annotations. On the other hand, these 11 benchmarks cover various complex domains and diverse scenes, ranging from first-person to third-person perspectives, and from indoor to outdoor environments. Hence, our MVBench is a preferable choice to evaluate the general capability of MLLMs for open-world temporal understanding. More importantly, these benchmarks provide the ground truth for MVBench which guarantees evaluation fairness and accuracy, avoiding biased scoring of LLMs [81, 49].
@@ 3 MVBench
In this section, we present our MVBench in detail. We first design the temporal tasks in Tab. 1, and then automatically generate multiple-choice QAs for evaluation in Fig. 2.
@@ 3.1 Temporal Task Definition
Action. (1) Action Sequence: Retrieve the events occurring before or after a specific action. (2) Action Prediction: Infer the subsequent events based on the current actions. (3) Action Antonym: Distinguish the correct action from two inversely ordered actions. (4) Fine-grained Action: Identify the accurate action from a range of similar options. (5) Unexpected Action: Detect surprising actions in videos characterized by humor, creativity, or magic. Object. (6) Object Existence: Determine the existence of a specific object during a particular event. (7) Object Interaction: Identify the object that participates in a particular event. (8) Object Shuffle: Locate the final position of an object in an occlusion game. Position. (9) Moving Direction: Ascertain the trajectory of a specific object’s movement. (10) Action Localization: Determine the time period when a certain action occurs. Scene. (11) Scene transition: Determine how the scene transitions in the video. Count. (12) Action Count: Calculate how many times a specific action has been performed. (13) Moving Count: Calculate how many objects have performed a certain action. Attribute. (14) Moving Attribute: Determine the appearance of a specific moving object at a given moment. (15) State Change: Determine whether the state of a certain object changes throughout the video. Pose. (16) Fine-grained Pose: Identify the accurate pose category from a range of similar options. Character. (17) Character Order: Determine the order in which the letters appear. Cognition. (18) Egocentric Navigation: Forecast the subsequent action, based on an agent’s current navigation instructions. (19) Episodic Reasoning: Perform reasoning on the characters, events, and objects within an episode of a TV series. (20) Counterfactual Inference: Consider what might happen if a certain event occurs.
@@ 3.2 Automatic QA Generation
With the guidance of temporal task definitions, we next collect and annotate videos for each task. Specifically, we design an automatic QA generation paradigm in Fig. 2, which efficiently converts open-sourced video annotations into multiple-choice QAs for evaluating MLLMs.
@@
QA Generation. Considering that not all the annotations of selected datasets follow the multiple-choice QA format, we automatically convert the video annotations into this format via LLMs. Specifically, we first use ChatGPT [54] to generate a question for each video, based on the task definition. Then, we create the corresponding answer options as follows. (1) Template-Based Construction: For most questions, we construct the option candidates directly from the ground truth annotations. For example, the candidates for the Action Antonym task contain the correct action, its opposite action, and a not-sure choice. In the case of the Moving Direction task, the option candidates consist of four directions (i.e., up, down, left, right) and the stationary state. (2) LLM-Based Generation: For the Unexpected Action task in particular, we leverage ChatGPT for converting open-ended QAs into multiple-choice QA with answer options. Note that, we use the multiple-choice format instead of the open-ended one, for evaluation correction and fairness. This is mainly because the open-ended answer has to be scored by LLMs or user studies, which may either introduce evaluation bias or manual intervention. Ultimately, we produce 200 multiple-choice QA pairs for each temporal understanding task. More details of QA generation for all the tasks can be found in the appendix.
@@ 4.2 Progressive Multi-Modal Training
| Random | - | 27.3 | 25.0 | 25.0 | 33.3 | 25.0 | 25.0 | 33.3 | 25.0 | 33.3 | 25.0 | 25.0 | 25.0 | 33.3 | 25.0 | 33.3 | 33.3 | 25.0 | 33.3 | 25.0 | 20.0 | 30.9 |
| Image MLLMs: Following [11], all models take 4 frames as input, with the output embeddings concatenated before feeding into the LLM. | ||||||||||||||||||||||
@@
| Video MLLMs: All models take 16 frames as input, with the exception of VideoChatGPT, which uses 100 frames. | ||||||||||||||||||||||
@@ 5.1 Results on MVBench
Additionally, we have included an evaluation of the powerful GPT-4V [55] on our MVBench. The results show that while GPT-4V [55] achieves satisfactory performance, particularly in demonstrating its capacity for temporal understanding, our VideoChat2 still manages to surpass it, increasing accuracy by 7.6%. This further underscores the superiority of our model in handling a broader range of tasks.
@@ 6 Conclusion
Appendix A Training Hyperparameters
The hyperparameters used in different stages of training are listed in Tab. 10. We adopt TSN [71] sampling for all the videos as previous methods [41, 72, 40]. For both Stage1 and Stage2, we employ large-scale image and video caption data, as outlined in the main manuscript. During Stage3, we make use of diverse instruction data and incorporate LoRA modules [24] into the LLM with a rank of 16, an alpha value of 32, and a dropout rate of 0.1. We apply flash attention [12] to expedite the training process.
@@ Appendix D Results on Challenging Video QA
In Tabs. 16 and 17, we extend the evaluation of our VideoChat2 to other challenging video benchmarks i.e., NExT-QA [80], STAR [78] and TVQA [34]. Different from the previous methods [90], which provide answers by comparing the likelihood of different options, we output the options directly, following the protocol of MVBench. Our results indicate that VideoChat2 not only holds its own against current SOTA methods [73, 90] on NExT-QA but also markedly outperforms them on STAR and TVQA. This underscores the effectiveness and robustness of VideoChat2.
@@
10cm
VideoChat2
58.5
2
[…]
10cm
VideoChat
42.5
3
[…]
14cm
LLaMA-Adapter
41.5
4
[…]
14cm
InstructBLIP
40.5
5
[…]
14cm
BLIP2
40.0
6
[…]
10cm
VideoChatGPT
39.5
7
[…]
14cm
LLaVA
38.5
8
[…]
10cm
VideoLLaMA
32.5
9
[…]
14cm
mPLUG-Owl-I
31.5
10
[…]
14cm
Otter-I
28.5
11
[…]
14cm
MiniGPT-4
8.0
@@
Appendix E Leaderboards and Analyses
To facilitate a clear comparison of different open-sourced MLLMs, we present the leaderboards for different tasks on MVBench in Tab. 18. Overall, our VideoChat2 achieves the highest rank across 15 tasks.
@@ Appendix E Leaderboards and Analyses
Appendix F Qualitative Results